OpenStudy (anonymous):

0+1+2+3+4 write the series with summation notation

OpenStudy (anonymous):

10

OpenStudy (anonymous):

\[\sum_{k=0}^4k\]

OpenStudy (anonymous):

sum n ( n +1)/2 = 4 * 5/ 2 = 10

OpenStudy (anonymous):

I came up with \[\sum_{i=1}^{5}i\]

OpenStudy (anonymous):

with the sum from i = 1 to 5, of i, you'll get 1 + 2 + 3 + 4 + 5